### v2

# PyraTrans: Attention-Enriched Pyramid Transformer for Malicious URL Detection

Abstract
Although advancements in machine learning have driven the development of malicious URL detection technology, current techniques still face significant challenges in their capacity to generalize and their resilience against evolving threats. In this paper, we propose PyraTrans, a novel method that integrates pretrained Transformers with pyramid feature learning to detect malicious URL. PyraTrans utilizes a pretrained CharBERT as its foundation and is augmented with three interconnected feature modules: 1) Encoder Feature Extraction, extracting multi-order feature matrices from each CharBERT encoder layer; 2) Multi-Scale Feature Learning, capturing local contextual insights at various scales and aggregating information across encoder layers; and 3) Spatial Pyramid Attention, focusing on regional-level attention to emphasize areas rich in expressive information. The proposed approach addresses the limitations of the Transformer in local feature learning and regional relational awareness, which are vital for capturing URL-specific word patterns, character combinations, or structural anomalies. In several challenging experimental scenarios, the proposed method has shown significant improvements in accuracy, generalization, and robustness in malicious URL detection. For instance, it achieved a peak F1-score improvement of 40% in class-imbalanced scenarios, and exceeded the best baseline result by 14.13% in accuracy in adversarial attack scenarios. Additionally, we conduct a case study where our method accurately identifies all 30 active malicious web pages, whereas two pior SOTA methods miss 4 and 7 malicious web pages respectively. Codes and data are available at: https://github.com/Alixyvtte/PyraTrans.
Index Terms:
Cybersecurity, Phishing Detection, Transformer, Multi-Scale Learning, Pyramid Attention
I Introduction
Malicious URLs, systematically engineered by cybercriminals for illicit activities such as scams, phishing, spam, and malware distribution, constitute a significant cybersecurity risk. These URLs directly threaten user and organizational security, leading to privacy breaches, data theft, extortion, and compromising the integrity of devices and networks. Vade’s 2023-Q3 report highlights a substantial escalation in phishing and malware incidents. The period saw malware volumes nearing the highest quarterly record, second only to Q4 2016’s 126.8 million instances. A notable 173% increase in phishing incidents was reported over the previous quarter, totaling 493.2 million, the highest for any Q3 since systematic tracking began in 2015 [1].
Predominantly, cyber attackers leverage deceptive hyperlinking as a key strategy in phishing schemes, often imitating credible entities like Microsoft, Google, and Facebook [2]. The ability to alter the display text of hyperlinks in HTML exacerbates the threat by camouflaging the true malicious nature of these URLs. This tactic poses a significant challenge to the detection of malicious URLs. Traditional database-centric detection platforms, such as Phishtank, reliant on community submissions, often face delays in identifying and mitigating new threats. The urgency for advanced systems capable of adapting to the dynamic cybersecurity environment and effectively identifying deceptive malicious URLs is highlighted by this lag.
Traditional methods for detecting malicious URLs, such as blacklist, heuristic, and rule-based approaches, face limitations as they require a understanding of URL structures and nature (as depicted in Figure 1[3, 4]. Blacklist detection encounters issues such as delayed updates and struggles with identifying unknown malicious URLs. Heuristic methods show performance variations across different malicious URL samples, while rule-based detection needs manual maintenance, making it challenging to address emerging threats and their variations[5, 6]. These limitations highlight the need for utilizing more advanced machine learning methods. On the other hand, earlier studies illustrates that malicious URLs display highly discernible string patterns, including length, the quantity of dots, and specific words [7, 8]. These patterns play a vital role in threat analysis and provide the groundwork for training sophisticated classifiers.
TABLE I: Example of the BERT token sequence extraction from amazon web page.
URL https://www.contactmailsupport.net/customer-service/amazon/
Token Sequence ’[CLS]’,’https’,’:’,’/’,’/’,’www’,’.’,’contact’,’##mail’,’ ##su’,’##pp’, ’##ort’,’.’,’net’,’/’,’customer’,’-’,’service’, ’/’,’am’, ’##az’,’##on’,’/’,’[SEP]’
Advancements in deep learning have significantly propelled the development of malicious URL detection systems [9]. Utilizing stacked nonlinear layers, deep learning [10, 11] learns to extract abstract, dense representations from raw initial feature vectors, such as character-level, word-level, or n-gram-based features, circumventing the need for intricate expert feature engineering. Convolutional Neural Networks (CNNs) have been the predominant deep learning models in prior malicious URL detection efforts like URLNet [12], TException [13], and Grambeddings [14], where CNNs capture local URL patterns to achieve SOTA performance. However, the technical limitations of CNNs make it increasingly difficult to achieve significant improvements in the development of CNN-based malicious URL detection models.
Recently, pretrained Transformer frameworks, such as BERT [15], have demonstrated promising prospects in various fields including natural language understanding, computer vision, and bioinformatics. This novel computational architecture and training paradigm offer new opportunities in malicious URL detection [16, 17]. They not only have stronger contextual learning capabilities but are also purely data-driven, end-to-end models. Table I presents an example of a BERT token sequence generated from a URL. However, despite the success of pretrained Transformer across multiple disciplines, their technological advantages have yet to be fully explored and manifested in malicious URL detection. Additionally, when applying Transformers to URLs, extra design considerations are essential for local feature learning and focused attention on key segments, as these aspects are crucial for capturing specific word patterns, character combinations, or structural anomalies in URLs.
Figure 1: Some major parts in a URL.
In this study, we propose PyraTrans, a novel appraoch developed on the foundation of pretrained Transformers, aimed at effectively detecting malicious URLs. This method exploits the various encoding layers of pre-trained Transformers and amalgamates them with multi-scale feature extraction and pyramid attention. Our approach, when tested across a broad range of public benchmarks datasets, exhibited superior performance over previous methods, proving to be more generalizable and robust, particularly in challenging situations like adversarial attacks, cross-dataset evaluation, class-imbalanced learning. We also conduct a case study to evaluate the model in real-world applications.
The main contributions of this paper are as follows:
  • Our method establishes new SOTA performance in malicious URL detection. Throughout extensive experimental evaluations, it consistently demonstrates superior generalization and robustness compared to previous methods. Additionally, it excels in learning under class imbalance and shows improved multi-classification performance, all while reducing the reliance on extensive training data.
  • This work introduces a straightforward multi-order URL feature generation approach by leveraging the multi-level encoding structure of Transformers. Additionally, we develop a multi-scale feature learning module to address the limitations of Transformers in extracting local string pattern from URLs.
  • We combine Transformers with spatial pyramid attention in our method, enabling it to assign regional-level attention. This allows our model to focus on segments with high information density in URLs while reducing the impact of noisy parts, such as random strings.
  • Our research reveals that even previously most popular methods, when subjected to comprehensive evaluations and practical case studies, can exhibit significant limitations. This underscores the critical need for robust case studies in assessing real-world applicability and for a more nuanced evaluation system that considers a method’s generalization and robustness.
The paper unfolds as follows: Section II conducts a literature review, while Section III outlines the datasets used. In Section IV, we provide a thorough explanation of the architecture and key components within our model. V details extensive experiments on malicious URL detection, benchmarking against baseline methods. A case study is then provided in Section V-E, and our conclusions are summarized in Section VII.
II Related Work
Malicious URL detection has a long-standing history in research. In this paper, we primarily review some recent works relevant to our study, which can be categorized into two types: CNN-based approaches and Transformer-based approaches.
II-A CNN-based Detection
Huang et al. [18] proposed a network that incorporates convolutional layers and two capsule network layers to learn the embedding representations of URLs. Wang et al. [19] combined CNNs and RNNs to extract key features for measuring content similarity, integrating these with static lexical features extracted from URLs using Word2Vec for their detection model. URLNet [12] introduced a dual-channel CNN approach for learning both character and word-level embeddings, combining these at the model’s top. This method not only achieved state-of-the-art performance at the time but also inspired numerous subsequent studies [13, 20, 21, 22], which all adopted the dual-channel feature extraction concept of URLNet. Recently, Bozkir et al. [14] developed Grambeddings, a neural network that effectively combines CNNs, LSTMs, and attention mechanisms. This network represents URL features through n-grams and has shown performance that surpasses URLNet in certain aspects.
The application of traditional neural networks in this field has seen widespread adoption. However, recent research suggests that their performance appears to have reached a plateau, leaving limited room for further improvement. Moreover, although these methods have advanced malicious URL detection, they still rely on manually initialized features at different levels (characters, words, or n-grams). In contrast, our proposed method is purely data-driven, requiring no manual engineering, and ingeniously implements feature extraction at both subword and character levels.
TABLE II: The statistical analysis of our datasets
Dataset Sample Sizes Avg Length Benign TLDs Malicious TLDs
malicious5 benign total malicious benign .com ccTLDs others .com ccTLDs others
GramBeddings1 400,000 400,000 800,000 86.24 46.38 52.17% 12.04% 35.79% 60.10% 11.82% 28.08%
Mendeley2 35,315 1,526,619 1,561,934 37.15 35.82 61.97% 0.93% 37.10% 72.86% 1.61% 25.53%
Kaggle 13 316,251 316,252 632,503 64.68 58.30 77.46% 0.63% 21.92% 50.59% 10.61% 38.8%
Kaggle 24 213,037 428,079 641,116 64.13 57.69 74.27% 6.61% 19.12% 46.62% 7.74% 45.65%
  • 1,2
    These are used for binary classification, download using Grambeddings and Mendeley links.
  • 3
    This is used for binary cross dataset test, download using this link.
  • 4
    This is used for multiple classification, download using this link.
  • 5
    Indicates malicious URLs in binary test and the total of malicious, defacement, and phishing URLs in multiple test.
II-B Transformer-based Detection
Chang et al. [11] fine-tuned a BERT model, initially pretrained on English text, using URL data for detecting malicious URLs. URLTran [10] comprehensively analyzed transformer models for phishing URL detection, demonstrating their feasibility and exploring various hyperparameter settings. However, these approaches, due to limited technical modifications, could not overcome the bias between the pretrained data domain and the task domain. The study in [23] employed a Transformer with a hybrid expert network for URL classification. Xu et al. [24] used a lightweight Transformer-based model. Although these methods achieved good performance, they did not fully leverage the advantages of pretraining. In the work of Wang et al. [25], a domain-specific BERT architecture was pretrained from scratch for URL applications. While this approach offers many benefits, it requires extensive URL data, substantial computational resources, and extensive training time.
III Large Scale URL Dataset
The four datasets used for training, validation and testing are publicly available. These datasets share a similar schema, consisting of the browsing URL and a corresponding label indicating whether the URL has been identified as malicious or benign. Upon analyzing their statistical data, including sample size, average URL length, and top-level domain (TLD) types, as shown in TableII, we discovered certain variations in the URL data across these datasets, which contribute to a more comprehensive evaluation of our model.
GramBeddings Dataset: As provided by Grambeddings[14], this dataset comprises 800,000 samples, equally divided into 400,000 malicious and 400,000 benign URLs. Malicious URLs were collected from websites such as PhishTank and OpenPhish, spanning the period from May 2019 to June 2021. Long-Term and Periodical Sampling, as well as Similarity Reduction techniques, were applied to the malicious data. The benign URLs were iteratively crawled from Alexa and the top 20 most popular websites in 20 different countries, and then randomly sampled. As a result, this dataset presents the highest diversity and sample size compared to others, while also demonstrating an equal number of instances per class. As shown in the table, the average length of malicious URLs is significantly longer than that of benign URLs, approaching twice the length, ensuring the similarity of the class-level average length distribution. In terms of top-level domain (TLD) features, because different domains are generally more difficult to obtain in malicious URLs, this dataset has improved the domain-level diversity of malicious URLs by setting a low ratio of unique domains to total domains, achieving a ratio similar to that of benign URLs, with .com at 60.10% and ccTLDs at 11.82%, whereas for benign URLs, it is .com at 52.17% and ccTLDs at 12.04%.
Mendeley Dataset: From Mendeley Data [26], this dataset consists of 1,561,934 samples, with a significant skew towards benign URLs (1,526,619) compared to 35,315 malicious URLs. The samples were crawled using the MalCrawler tool and validated using the Google Safe Browsing API[27]. This dataset exhibits a notable class imbalance at a ratio of approximately 1:43. In terms of average length, both malicious and benign URLs demonstrate similar values, and the diversity of top-level domains (TLDs) is limited, primarily concentrated in .com, accounting for 61.97% and 72.86%, respectively, with ccTLDs representing only 0.93% and 1.61%. Although this may pose a risk of misleading the training processes by capturing inadequate syntactical or semantic features, considering the potential encounter with such data distribution in real-world scenarios, we chose to employ this dataset for model evaluation.
Kaggle Dataset: The Kaggle 1 and Kaggle 2 datasets are both derived from the Kaggle website. Kaggle 1 is designed for binary classification experiments, while Kaggle 2 is intended for multi-classification tasks. Kaggle 1 consists of 632,503 samples, evenly distributed between malicious and benign URLs. In comparison to the Mendeley dataset, this dataset demonstrates disparity in the average length of malicious and benign URLs, with the samples presenting a more balanced composition between the two classes. Notably, within the malicious samples of this dataset, there is a noticeable higher ratio of ccTLDs in TLDs compared to the benign URLs, accounting for 10.61%, while benign URLs only make up 0.63%. Additionally, the proportion of .com domains is 50.59% for malicious URLs, whereas it is 77.46% for benign URLs.
The Kaggle 2 dataset consists of four classes: benign (428,079), defacement (95,306), phishing (94,086), and malicious (23,645). The benign class contains positive samples, while the other three classes contain negative samples of different types. We observe that the .com TLDs are dominant in benign URLs, accounting for 74.27% of the total. The ccTLDs are slightly more frequent in this dataset (6.61%) than in the other two datasets, while the other gTLDs represent 19.12% of the benign URLs. For the negative samples, the .com TLDs are less prevalent, with a frequency of 46.62% across all three classes. The ccTLDs and other gTLDs have higher frequencies of 7.74% and 45.65%, respectively, in the negative samples than in the benign ones.
The distinctive composition and Top-Level Domain (TLD) distribution in each dataset offer a comprehensive foundation for assessing the efficacy of our method across diverse web domains. This enables robust testing under different real-world scenarios.
IV Methodology
The proposed method employs the CharBERT network structure as the backbone network, integrating an encoder feature extraction module, a multi-scale learning module, and a Spatial Pyramid Attention module.
The overall model structure is depicted in Fig. 3.
  • Backbone Network: The CharBERT model improves URL data interpretation and analysis with its advanced subword and character-level embedding which extends based on BERT.
  • Encoder Feature Extraction: This module extracts multi-layer encoder features from CharBERT, aiming to capture URL representations ranging from low-level to high-level.
  • Multi-scale Learning: The module conducts multi-scale local information extraction from multi-layer encoder features and captures the relational information between different encoder feature layers.
  • Spatial Pyramid Attention: This module differentially weights different regions of the feature, highlighting local spatial correlations, allowing flexible focus on information-rich segments in URLs. This contrasts with Transformer’s Multi-Head Attention, which prioritizes positional relationships.
Figure 2: The architectural diagram of BiGRU.
Figure 3: PyraTrans: Composed of Four Core Components. CharBERT, the backbone network for learning character and subword-level features; Encoder Feature Extractor, Multi-Scale Feature Learning, and Spatial Pyramid Attention module for acquiring multi-order, multi-scale, attention-weighted features.
IV-A Backbone Network
We utilize CharBERT[28], an extension of BERT integrating the Transformer architecture with a dual-channel framework, as our pretrained backbone network to capture both subword and character-level features. CharBERT capture character-level information in token sequences through two modules: (1) the Character Embedding Module, encoding character sequences from input tokens and (2) the Heterogeneous Interaction Module, combines features from both character and subword channels, and then independently separates them into distinct representations as input for the encoder layer.
The character-aware embedding of each token is primarily generated through two components: the encoding of individual characters and subword units. These two components are integrated via a dual-channel architecture. To establish contextual character embeddings, we utilize a bidirectional Gated Recurrent Unit (BiGRU) layer. The BiGRU employs a bidirectional recurrent neural network with only the input and forget gates[29]. The architecture diagram of the BiGRU is depicted in Figure 2.
Assuming $x$ denotes the input data, and $h$ represents the output of GRU unit. $r$ is the reset gate, and $z$ is the update gate. $r$ and $z$ decide how to get the new hidden state $h_{t}$ from the previous hidden state $h_{t-1}$ calculation. The update gate controls both the current input $x_{t}$ and the previous memory $h_{t-1}$, and outputs a numerical value $z_{t}$ between 0 and 1. The calculation formula is as follows:
$z_{t}=\sigma(W_{z}[h_{t-1},x_{t}]+b_{x})$ (1)
where $z_{t}$ determines the extent to which $h_{t-1}$ should influence the next state, $\sigma$ is the sigmoid activation function, $W_{z}$ is the update gate weight, and $b_{z}$ is the bias. The reset gate regulates the influence of the previous memory $h_{t-1}$ on the current memory $h_{t}$, removing it if deemed irrelevant.
$r_{t}=\sigma(W_{r}[h_{t-1},x_{t}]+b_{x})$ (2)
Then creating new memory information $h_{t}$ using the update gate:
$\widetilde{h_{t}}=tanh(W_{h}[r_{t}h_{t-1},x_{t}]+b_{h})$ (3)
The output at the current moment can be obtained:
${h_{t}}=(1-z_{t})h_{t-1}+z_{t}\widetilde{h_{t}}$ (4)
The current hidden layer state of the BiGRU is influenced by the current input $x_{t}$, the forward hidden state $\overset{\xrightarrow{}}{h_{t-1}}$, and the output $\overset{\xleftarrow{}}{h_{t}}$ of the reverse hidden layer state:
$\overset{\xrightarrow{}}{h_{t}}=\overset{\xrightarrow{\hskip 14.22636pt}}{GRU}(\overset{\xrightarrow{\hskip 8.5359pt}}{h_{t-1}},x_{t})(t=1,2,…,d)$ (5)
$\overset{\xleftarrow{}}{h_{t}}=\overset{\xleftarrow{\hskip 14.22636pt}}{GRU}(\overset{\xleftarrow{\hskip 8.5359pt}}{h_{t+1}},x_{t})(t=d,d-1,..,1)$ (6)
$h_{t}=w_{t}\overset{\xrightarrow{}}{h_{t}}+v_{t}\overset{\xleftarrow{}}{h_{t}}+b_{t}=BiGRU(x_{t})$ (7)
The $GRU$ represents the nonlinear transformation of the input, incorporating the degradation indicator into the associated GRU hidden state. $w_{t}$ and $v_{t}$ denote the weights of the forward hidden layer state $\overset{\xrightarrow{}}{h_{t}}$ and reverse hidden state output $\overset{\xleftarrow{}}{h_{t}}$ of the bidirectional GRU at time $t$, respectively, $b_{t}$ represents the bias corresponding to the hidden state at time $t$.
In the generation of character embeddings, we represent an input sequence as ${w_{1},...,w_{i},...,w_{m}}$,, where $w_{i}$ is a subword tokenized using Byte Pair Encoding (BPE), and $m$ is the length of the sequence at the subword level. Each token $w_{i}$ consists of characters ${c^{i}_{1},...,c^{i}_{n_{i}}}$, where $n_{i}$ represents the length of the subword. The total character-level input length is denoted as $N=\sum\limits_{i=1}^{m}n_{i}$, where $m$ is the number of tokens. The formulation of the processing is as follows:
$e^{i}_{j}=W_{c}\cdot c^{i}_{j};\;h^{i}_{j}=BiGRU(e^{i}_{j});$ (8)
Here, $W_{c}$ is the character embedding matrix, and $h^{i}_{j}$ denotes the representation of the $j$-th character within the $i$-th token. The BiGRU processes characters across the entire input sequence of length $N$ to generate token-level embeddings. Then connect the hidden states of the first and last characters in each token, as follows:
$h_{i}(x)=[h^{i}_{1}(x);h^{i}_{n_{i}}(x)]$ (9)
Let $n_{i}$ be the length of the $i$-th token, and $h_{i}(x)$ be the token-level embedding from characters, enabling contextual character embeddings to capture complete word information.
Figure 4: The architecture of Heterogeneous Interaction Module.
The heterogeneous interaction module fuses and separates the token and character representations after each transformer layer. The structure shown in Fig4. This module uses different fully-connected layers to transform the representations, and then concatenates and integrates them by using a CNN layer, as follows:
$t^{{}^{\prime}}_{i}(x)=W_{1}*t_{i}(x)+b;\;h^{{}^{\prime}}_{i}(x)=W_{2}*h_{i}(x)+b_{2}$ (10)
$w_{i}(x)=[t^{{}^{\prime}}_{i}(x);h^{{}^{\prime}}_{i}(x)];\;m_{j,t}=tanh(W^{j}_{3*w_{t:t+s_{j}-1}}+b^{j}_{3})$ (11)
where $t_{i}(x)$ is the token representation, $W$ and $b$ are the parameters, $w_{t:t+s_{j}-1}$ is the concatenation of the embeddings of $(w_{t},...,w_{t+s_{j}-1})$, $s_{j}$ is the window size of the $j$-th filter, and $m$ is the fused representation, which has the same dimension as the number of filters.
Next is a fully connected layer with GELU activation[30], used to restore the fused features to two channels, with residual connection added to preserve the original information of each channel.
$m^{t}_{i}(x)=\Delta(W_{4}*m_{i}(x)+b_{4});\;m^{h}_{i}(x)=\Delta(W_{5}*m_{i}(x)+b_{5})$ (12)
$T_{i}(x)=t_{i}(x)+m^{t}_{i}(x);\;H_{i}(x)=h_{i}(x)+m^{h}_{i}(x)$ (13)
$\Delta$ is the activation function GELU, and T and H as the representations of the two channels. After the residual connection, a layer normalization operation is applyed. The fusion and separation process can enrich the mutual representations of the two channels, while preserving the specific features of the tokens and characters. The pre-training tasks can also enhance the differentiation of the dual-channel framework.
IV-B Encoder Feature Extraction
Pre-trained language models such as BERT use multiple layers of Transformer encoders to learn semantic knowledge from large-scale corpora, and then fine-tune them for specific downstream tasks. Most BERT-based classification models depend on the [CLS] feature of the final layer, which summarizes the semantic information of the whole input sequence. However, Jawahar et al. [31] show that BERT can learn various information across layers, such as phrase-level details in lower layers, syntactic information in middle layers, and rich semantic features in higher layers. They apply k-means clustering to the BERT layer representations and measure the cluster quality by using Normalized Mutual Information (NMI). They find that lower BERT layers are better at encoding phrase-level information, as indicated by higher NMI scores, as shown in Table III. Deeper encoder layers are more effective in handling long-range dependency information.
Although each layer in the BERT family of models takes the output features of the previous layer as input for computation, multiple intricate calculations within each layer’s processing may still result in potential degradation of lower-level and mid-level features, which is detrimental to the complete feature learning process. Li et al. [32] use feature concatenation to integrate aspect features from each layer of BERT for aspect term sentiment classification. This approach, instead of relying only on the final layer for classification features, effectively enhances classification performance by leveraging the distinct features learned at each layer of BERT.
TABLE III: Performance of span representation clustering derived from various layers of BERT[31]
Layer 1 2 3 4 5 6
NMI 0.38 0.37 0.35 0.3 0.24 0.2
Layer 7 8 9 10 11 12
NMI 0.19 0.16 0.17 0.18 0.16 0.19
Similar to prior research, we extract outputs from each encoding layer in CharBERT. However, instead of concatenating these layer-wise features, we reorganize them into a higher-dimensional matrix. In this restructured feature matrix, layers function akin to channels in an image. The feature process is as follows: Consider a sequence of outputs $k_{1},k_{2},...,k_{n}$ and $u_{1},u_{2},...,u_{m}$, where each output $k_{i}$ and $u_{j}$ has a rank of $(H,W,C)$, representing the outputs of CharBERT’s word-level and character-level encoders at various layers. $H$ is the batch size, $W$ is the fixed URL sequence length (200 in our model), and $C$ is a 768-dimensional vector for each merged hidden layer output in CharBERT. For example, $k_{1}$ and $u_{2}$ be two tensors representing the sequence and character embeddings, respectively. Let $w$ be the sequence length, and $d$ be the embedding dimension:
$k_{1}$ $=\begin{bmatrix}x_{11}^{1}&x_{12}^{1}&\cdots&x_{1d}^{1}\\ x_{21}^{1}&x_{22}^{1}&\cdots&x_{2d}^{1}\\ \vdots&\vdots&\ddots&\vdots\\ x_{w1}^{1}&x_{w2}^{1}&\cdots&x_{wd}^{1}\end{bmatrix},\quad u_{1}$ $=\begin{bmatrix}x_{11}^{2}&x_{12}^{2}&\cdots&x_{1d}^{2}\\ x_{21}^{2}&x_{22}^{2}&\cdots&x_{2d}^{2}\\ \vdots&\vdots&\ddots&\vdots\\ x_{w1}^{2}&x_{w2}^{2}&\cdots&x_{wd}^{2}\end{bmatrix}.$ (14)
Afterwards, we use one-dimensional convolution to fuse the concatenated channel features, reducing their dimensionality to the original values of each channel. Here, $C$ represents concatenation, $K_{fuse}$ denotes convolution, and $Y$ is the resulting fused tensor:
$Y=K_{fuse}(C(k_{1},u_{1})^{T})=\begin{bmatrix}y_{11}&y_{21}&\cdots&y_{m1}\\ y_{12}&y_{22}&\cdots&y_{m2}\\ \vdots&\vdots&\ddots&\vdots\\ y_{1d}&y_{2d}&\cdots&y_{md}\end{bmatrix}$ (15)
By stacking the merged output along the new dimension 0, we form a tensor $F$ of rank $(N,H,W,C)$, where $N$ is the number of layers (12 in our model). To align the multilevel features for subsequent analytical tasks, the tensor elements were rearranged by permuting dimensions 0 and 1. This resulted in a tensor $F^{\prime}=(H,N,W,C)$, which served as the stacked feature input for the next multi-attention module.
IV-C Multi-scale Learning
In extracting URL embedding features, the standard architecture of the Transformer model lacks specialized design for capturing local features. This means it can comprehend the context of the entire input sequence but may not focus on local details. However, capturing local features in URLs is crucial, especially in applications like security analysis, fraud detection, or content categorization. Local features, including specific word patterns, character combinations, or structural anomalies, can be indicative of the nature and intent of a URL.
To more effectively capture these local features, we augment our architecture with multi-scale feature learning, specifically designed for local feature extraction. This module is based on depthwise separable convolutions (DSConv) [33], offering reduced floating-point operations and enhanced computational efficiency. We employ dilated convolutions with varying dilation rates to capture multi-scale information, serving as fundamental operators for expanding the network’s depth and breadth. Formally, the high-dimensional feature represented by outputs from multiple encoder layers is denoted as $M\in\mathbb{R}^{C\times H\times W}$, where $C$ is the number of channels, $H$ is the height, and $W$ is the width. The process begins by applying a single DSConv (conv3 x 3) to $M$, extracting common information denoted as $F_{0}$ for each branch. Specifically:
$F_{0}=K_{0}(M)$ (16)
$K_{0}$ denotes a depthwise separable conv3 x 3 operation, and dilated DSConv3x3 with different rates are applied to $F_{0}$ across branches ($K_{i}$ for branch $i$, $N$ branches). Contextual information from multiple scales is integrated through element-wise summation using a residual connection, termed:
$F_{i}=K_{i}(F_{0}),i=1,2,...,N$ (17)
$F=\sum_{i=0}^{N}F_{i}$ (18)
Since the concatenation operation substantially amplifies channel count, leading to increased computational complexity and network parameters. Thus, we opt for element-wise summation. Lastly, aggregated features are reshaped using a 1 × 1 standard convolution. Formally:
$Q=K_{fuse}(F)+M$ (19)
$K_{fuse}$ denotes the standard conv1 x 1 operation for fusing additional information at different scales. The original feature map $M$ is integrated as a residual connection [34], aiding gradient flow and facilitating effective training. In our experiments, dilation rates of $[1,2,4,8]$ are utilized to capture contextual information at various scales.
Within the multi-scale learning module, a straightforward element-wise summation of features from various scales may inadvertently diminish the importance of informative branches while according equal significance to all scales. To mitigate this problem, we employ a spatial pyramid attention mechanism[35], which effectively assesses subfields across multiple scales and adjusts branch weights, enhancing the overall performance.
IV-D Spatial Pyramid Attention
In the multi-scale learning module, a simple element-wise summation of features from different scales may inadvertently downplay the importance of informative branches, treating all scales equally. Additionally, while Transformers offer token-level attention, for URL feature learning, regional-level attention is crucial due to the presence of distinct information-dense areas (like domain names) and regions with noise (such as random parameters) in URLs. To address these issue, we integrate a Spatial Pyramid Attention module following our multi-scale learning [35].
The spatial pyramid attention mechanism comprises three key elements: point-wise convolution, spatial pyramid structure, and a multi-layer perceptron. The point-wise convolution aligns channel dimensions and consolidates channel information. The spatial pyramid structure incorporates adaptive average pooling of three different sizes, promoting structural regularization and information integration along the attention path. Multi-layer perceptron then extracts an attention map from the output of the spatial pyramid structure.
To be sepecific, we denoted adaptive average pooling and fully connected layer as $P$ and $F_{fc}$ respectively. The concatenation operation is represented as $C$, $\sigma$ denotes the Sigmoid activation function, while $R$ is referred to as resizing a tensor to a vector. The fused feature map after the multi-scale learning moule can be denoted as $Q\in\mathbb{R}^{C\times H\times W}$, the attention mechanism learns attention weights from the input and multiplies each channel in it by learnable weights to produce an output. The output of the spatial pyramid structure $S(Q)$ can be presented as:
$S(Q)=C(R(P(Q,4)),R(P(Q,2)),R(P(R,1)))$ (20)
Omitting the batch normalization and activation layers for the sake of clarity, the fundamental transformation $\sigma$ can be expressed as:
$\zeta(Q)=\sigma(F_{fc}(F_{fc}(S(Q))))$ (21)
In our experiments, the channel number $C$ is 12 according to the former process and we adopt 3-level pyramid average pooling. In the concluding phase of our network, we apply Mean Pooling to the weighted feature map along the fixed sequence length dimension. This outcome is then integrated with a dropout layer, followed by a fully connected layer that converts URL features into a binary class representation for prediction.
TABLE IV: Detection results of our proposed method vary with the number of layers employed
# layers(count) Accuracy Precision Recall F1-score AUC
2 0.9772 0.9811 0.9730 0.9771 0.9959
3 0.9837 0.9873 0.9799 0.9863 0.9963
4 0.9856 0.9917 0.9792 0.9854 0.9938
5 0.9860 0.9861 0.9858 0.9859 0.9998
12 0.9915 0.9949 0.9880 0.9914 0.9965
V Experiments
This section presents the detailed experimental setup and results to assess the effectiveness of our proposed method and compare it with baselines. Our experiments are primarily divided into the following components:
Figure 5: Detection results of selected baseline methods and our proposed method on GramBeddings dataset.
  1. 1.
    Assess stacked feature layer configurations for multi-layer extraction.
  2. 2.
    Explore data scale dependency with varied training dataset sizes.
  3. 3.
    Evaluate model generalization across datasets.
  4. 4.
    Multi-class classification.
  5. 5.
    Test robustness using adversarial samples.
  6. 6.
    Evaluate practicality with recent active malicious URLs.
Setup. The pre-trained CharBERT was trained on the English Wikipedia corpus, consisting of a total of 12GB and approximately 2,500 million words. Hyperparameter tuning during fine-tuning led to a batch size of 64, AdamW optimizer (initial learning rate: 2e-5, weight decay: 1e-4), 0.1 dropout rate, and 5 training epochs. We utilized PyTorch 2.0, NVIDIA CUDA 11.8, and Python 3.8, conducting training on NVIDIA A100 GPUs. The final model for each experiment was selected based on the best validation loss.
Baselines. In our experimental comparison, we chose the state-of-the-art models, URLNet and Grambeddings, as benchmarks for evaluating our proposed method. To ensure fairness and reproducibility, we obtained their code from GitHub repositories without making any modifications to the structure or hyperparameters, and applied them to our dataset. Specifically, for URLNet, we selected Embedding Mode 5, the most complex mode, as it exhibited superior performance in their original study.
V-A Evaluate the Effectiveness of Multi-level Features
We develop a multi-layer feature extraction module to distill semantic features from different layers of the Transformer encoders in our backbone network. Our approach involves stacking embedding outputs from different layers to form a multi-layered feature matrix, which we optimize to enhance URL semantic feature representation. Through this, we investigate the complexity of feature representation across the network and aim to demonstrate the effectiveness of multi-layered features in improving overall model performance.
We create a training corpus from the GramBeddings dataset, randomly selecting 128k URLs from the 80k available in the training set while maintaining the proportion of malicious and benign URLs. We also randomly sample 32k URLs for testing and validation respectively. The performance of various stacked configurations and the incremental gains are shown in Table IV. We observe a consistent improvement in evaluation metrics as we incrementally incorporate additional layers, namely the last 2, 3, 4, and 5 layers. Ultimately, the stacking of 12 embedding output layers achieves the best performance in URL detection tasks, demonstrating the effectiveness of integrating both lower and deeper layers in the model architecture.
V-B Comparison with Baselines
In this section, we compare the performance of PyraTrans with baselines using binary and multi-class malicious URL detection.
TABLE V: Detection results of selected baseline methods and our proposed method on Mendeley dataset.
Training Size Method Accuracy Precision Recall F1-score AUC
629,184 (60%) URLNet 0.9858 0.9475 0.3889 0.5515 0.9046
GramBeddings 0.9801 0.6137 0.3026 0.4053 0.8205
Our 0.9886 0.9104 0.5419 0.7027 0.9438
419,456 (40%) URLNet 0.9842 0.9810 0.3019 0.4617 0.8992
GramBeddings 0.9794 0.9984 0.0817 0.1510 0.8750
Our 0.9882 0.8998 0.5307 0.6677 0.9370
209,064 (20%) URLNet 0.9785 0.9653 0.0450 0.0860 0.7762
GramBeddings 0.9804 0.9677 0.1306 0.2301 0.7869
Our 0.9879 0.9131 0.5055 0.6507 0.9322
104,832 (10%) URLNet 0.9789 0.8382 0.0735 0.1351 0.7584
GramBeddings 0.9757 0.3879 0.1423 0.2082 0.8153
Our 0.9865 0.8540 0.4774 0.6125 0.9161
104,832 (5%) URLNet 0.9776 0.0000 0.0000 0.0000 0.6746
GramBeddings 0.9782 0.5647 0.1139 0.1896 0.7752
Our 0.9861 0.8706 0.4397 0.5843 0.8994
10,432 (1%) URLNet 0.9776 0.0000 0.0000 0.0000 0.4419
GramBeddings 0.9722 0.1808 0.0682 0.0991 0.6185
Our 0.9834 0.7632 0.3717 0.5000 0.8550
TABLE VI: Cross-dataset performance generalization
Cross-dataset Method Accuracy Precision Recall F1-score AUC
Gram/Kaggle URLNet 0.8823 0.8947 0.8666 0.8804 0.9492
Grambeddings 0.5214 0.5120 0.8595 0.6552 0.4647
Our 0.9138 0.9576 0.8641 0.9085 0.9705
V-B1 Binary classification
In the binary classification detection task, we used two datasets with significant differences. The first is the Grambeddings dataset, characterized by a balanced distribution of positive and negative samples and high diversity. The second is the Mendeley dataset, which exhibits extreme class imbalance and lower diversity. These datasets are used for evaluations in different detection scenarios. We explored the model’s dependency on training data size by varying it, starting from as low as 1%. Specifically, the experimented training sizes include 1% , 5%, 10% , 20%, 40% , 60% , and 80% . And the trained models are tested across all the test datasets.
Results on Grambeddings dataset: As shown in Figure 5, our proposed method achieves superior performance over the baseline method on the Grambeddings dataset, regardless of the size of the training set. Remarkably, our model demonstrates high proficiency even with scarce training data.
It is worth noting that, our method achieves remarkable performance with only 6,400 URLs (1%) for training, attaining an accuracy of 0.9358, which surpasses the baseline methods that range from 0.8349 to 0.8509. Furthermore, our model exhibits a high sensitivity in detection, with a recall of 0.9384, compared to the baseline recall of 0.7131 and 0.8254. The maximum gap in F1 score reached 0.1084.
Despite the gradual improvement in the baseline model’s performance with larger training samples, narrowing the gap with our method, our approach consistently achieves an accuracy of 0.9825 and an F1 score of 0.9824 using 80% of the training dataset. Our model outperformed the best baseline model with an accuracy and F1 score 0.0089 and 0.0091 higher. Although the difference seems small, it becomes significant when dealing with large-scale datasets. In conclusion, our model exhibits superior performance in accurately detecting malicious URLs on a balanced dataset, even with a small training set size.
Results on Mendeley dataset: To evaluate our method in real-world internet scenarios, where phishing sites are significantly outnumbered by legitimate web pages, we use the Mendeley dataset for further testing. This dataset contains 1,561,934 URLs, with a high imbalance ratio of about 43 to 1 between benign and malicious samples. This extreme imbalance poses a notable challenge to model performance, as it may cause a bias towards the abundant benign URL samples and increase the false positive rate when detecting malicious samples.
As shown in Table V, our method exhibits significant advantages in class-imbalanced scenarios compared to other approaches. With just 1% of training data, our model achieves an accuracy of 0.9837, with a Precision of 0.7632, surpassing the best baseline Precision of 0.1808. The F1 score of our method is four times higher than the best baseline performance. As training data increases, our F1 score reaches 0.7027, substantially exceeding the baseline peak of 0.5515, with a Recall 15.3 higher than the best baseline result. These experiments demonstrate the substantial improvements our method brings in identifying malicious URL samples, significantly reducing false positives. Moreover, the high AUC reflects our model’s confidence in the identified samples, indicating its accurate capture of key differences between malicious and benign samples.
We observe that while URLNet achieves high accuracy on larger datasets, it suffers from a high false negative rate, particularly on smaller datasets. Notably, URLNet fails to identify any malicious URLs when the training data size is reduced to below 5%. In contrast, Grambeddings shows significant sensitivity to data, with its performance in detecting malicious URLs varying greatly across different training sample sizes. For instance, at 60% training data, its Precision drops to 0.6137, and at 40%, its Recall falls to just 0.0817. Compared to these, our method demonstrates consistent and reliable performance across various data scales, indicating its robustness to small-scale and class-imbalanced data scenarios.
V-B2 Multi-classification
To evaluate our method in the context of complex cyber threats, we conduct a multi-class classification experiment, using a Kaggle 2 dataset [36] with four URL categories: benign (428,079), defacement (95,306), phishing (94,086), and malicious (23,645). Fig. 6 shows the results of our model and the baseline methods.
Fig.6 illustrates the performance of our method and the baseline methods across all four categories. Our method surpasses the baseline methods in each metric. The average ROC curve highlights our method’s efficacy with a TPR of almost 90% at a low FPR of 0.001, surpassing other methods that achieve around 75%. Grambeddings struggled in recognizing negative samples from various categories, resulting in a 50% F1 score for defacement and phishing URLs and an overall accuracy of 83.91%. URLNet achieved an overall accuracy of 97.07%, falling short of our model’s 98.57%. These results demonstrate the robustness and effectiveness of our method in complex multi-class classification tasks, indicating its potential as a promising solution for malicious URL detection in cybersecurity.
V-C Cross-dataset Testing
To evaluate the generalization of models and to identify any potential weaknesses or biases, we set up a cross-dataset testing experiment. We do this by training the model on the Grambeddings dataset and subsequently testing it on the Kaggle binary classification dataset. It is noteworthy that the Grambeddings and Kaggle datasets significantly differ in their data collection times and sources.
The results, as shown in Table VI, indicate a marked decline in performance of baseline methods on data not included in their training set, with URLNet’s accuracy dropping to 0.8823 and Grambeddings’ accuracy reducing to 0.5214 and 0.4647, respectively. In contrast, our model maintained high accuracy on external datasets, achieving an AUC of 0.9705. This underscores that the knowledge gained from one dataset by our method can be effectively generalized to others, even if the data was collected much later than our model’s pretraining period. This demonstrates the adaptability and long-term applicability of our approach to different data environments.
Figure 6: Detection results of selected baseline methods and our proposed method on multiple classification dataset.
V-D Evaluation against Adversarial Attacks
Cybercriminals employ adversarial attacks to bypass systems by exposing them to inaccurate, unrepresentative, or malicious data. We employed a Compound Attack technique as our threat model, which involves inserting an evasion character to a benign URL sample to create a real-world compatible malicious URL. This technique was first proposed by Maneriker et al.[10] and later applied and extended by Grambeddings[14].
The generation of adversarial samples entails the utilization of XLM-RoBERTa [37] for domain tagging in provided URLs. This process ensures a minimal tag count, involves the random insertion of hyphens in split parts, and includes the substitution of benign domains with malicious ones, resulting in the creation of an adversarial list.
To construct the AdvTest set, we merged 80K legitimate URLs and randomly sampled 40K malicious samples from the original validation data. Furthermore, we introduced 40K adversarial samples generated from benign URLs. This novel dataset presents a substantial challenge to the robustness of the model. Then we evaluate our and baseline models using this novel dataset.
As illustrated in Figure 7 and Table VII, baseline methods experience a significant decline in performance under adversarial sample attacks. The accuracy of URLNet decreases to 76.10%, and that of Grambeddings to 70.18%, while our model maintains an accuracy above 90%, with an AUC of 98.39%, exceeding URLNet by about 20%. At a fixed FPR of 0.01, our method achieves a TPR of nearly 75%, more than double the TPR of baseline methods, both under 30%. These results indicate the robustness of our approach to adversarial sample attacks, suggesting increased effectiveness in preventing malicious attack evasion in real-world scenarios.
Figure 7: Area under ROC curve under adversarial attack.
TABLE VII: Performance under adversarial attack.
Method ACC P R F1 AUC
URLNet 0.7610 0.8745 0.5635 0.6854 0.8040
Gram 0.7018 0.6137 0.8564 0.8564 0.8564
Our 0.9023 0.9738 0.8104 0.8846 0.9839
V-E Case Study
We conduct a series of case studies applying our detection model to active malicious web pages to evaluate its practical utility. In November 2023, we crawled 30 active phishing URLs reported and verified on PhishTank and tested them with our model trained on 30% of the Grambeddings dataset. For comparison, tests were also conducted with the best-performing URLNet and Grambeddings models, trained on the same dataset. Results indicate that our model detected all malicious URLs with 100% accuracy, while Grambeddings misclassified 7 URLs, resulting in 76% accuracy, and URLNet misclassified 4 URLs, with an accuracy of 86%. Table VIII lists URLs that were incorrectly classified, in order to provide a detailed perspective on the performance of each model.
URLNet failed in detecting four malicious URLs, which had relatively short strings. Given that benign URLs on the internet are typically simple, URLNet might have mistakenly classified these short malicious URLs as benign due to their length similarity. This suggests that URLNet relies excessively on URL string length for classification, demonstrating limited capability in recognizing semantics and specific patterns in real-life scenarios. Conversely, Grambeddings, which combines convolutional neural networks, long short-term memory networks, and attention layers, is a more complex system. It misclassified instances of both longer and shorter URLs, indicating a minor influence of URL length. However, the significant diversity among the seven misclassified malicious URLs implies that Grambeddings’ performance could be affected by various factors, and its learning system might not have developed sufficiently generalized discriminative patterns.
In contrast, only our method demonstrated consistent or even improved performance in real-world applications, showing its adaptability to various forms of malicious URLs. Analyzing the differences at a technical level, we attribute our approach’s distinct advantage over others to its comprehensive feature consideration,includeing character-aware token embeddings, multi-level and multi-scale feature learning, and regional-level attention.
VI Discussion
The proposed method, validated through a comprehensive set of experiments, has demonstrated robust, accurate, and reliable performance. Here, we briefly discuss the beneficial advantages and insights brought about by our proposed approach.
TABLE VIII: Cross-dataset performance generalization
Malicious url URLNet Grambeddings Our
https://bafybeibfyqcvrjmwlpipqkdyt2xr46cea7ldciglcbybfwtk7cieugcj3e.ipfs.infura-ipfs.io
http://798406.selcdn.ru/webmailprimeonline/index.html
https://79efc264-a0d7-4661-900b-a8bc1443be89.id.repl.co/biptoken.html
http://ighji.duckdns.org
https://www.minorpoint.lqoipum.top/
https://colstrues.com/s/jsrj
https://innovativelogixhub.firebaseapp.com/
https://sites.google.com/view/dejoelinoctskxo2bb
https://gtly.to/-H0PPiKyq
  • Note: We use the symbols ✓and ✗ to denote the correct and incorrect classification results, respectively.
  1. 1.
    End-to-End Architecture: ur proposed method is an end-to-end network utilizing pretrained CharBERT, which requires no manual feature initialization. It directly processes raw URLs and generates character-aware subword token embeddings. In contrast, previous studies typically necessitated manual initialization of character and word-level representations and relied on dual-path neural networks. Our approach streamlines the processing workflow while maintaining efficient feature extraction capabilities.
  2. 2.
    Evaluation Metrics: Our experiments show that the proposed method significantly improves accuracy, robustness, and generalizability, consistently delivering stable and effective detection across various testing scenarios. Meanwhile, leading baseline methods, such as URLNet and Gramembedings, displayed clear weaknesses: URLNet struggled in class-imbalanced scenarios with small datasets, while Gramembedings faced significant performance fluctuation. Furthermore, these methods varied in their effectiveness in generalization and adversarial robustness tests. This highlights the need for a comprehensive performance evaluation system that goes beyond specific experimental setups.
  3. 3.
    Case Studies: Previous research often overlooked the importance of case studies, but our work emphasizes the necessity of applying models directly to active malicious links to accurately reveal their real-world performance. Our case studies have shown that even methods excelling in experimental settings can face significant challenges in practical applications. Additionally, case studies offer an opportunity to thoroughly analyze a model’s feature learning capabilities and shortcomings in information capture patterns, allowing for a more comprehensive assessment of the model’s technical design.
VII Conclusion
We have proposed a novel transformer-based and pyramid feature learning system called PyraTrans for malicious URL detection. Our method effectively leverages knowledge transfer from pretrained models to URL contexts, dynamically integrates character and subword-level features, and incorporates three closely integrated feature learning modules for URL feature extraction. The key contributions of our approach are: 1) enabling end-to-end learning from raw URL strings without manual preprocessing; 2) adopting an interactive subword and character-level feature learning network architecture for improved character-aware subword representations; 3) conducting effective multi-level and multi-scale URL feature learning based on our proposed lightweight feature learning modules, addressing inherent limitations of the Transformer in local feature extraction and spatial awareness. We conduct extensive experiments on various URL datasets, demonstrating that our method consistently outperforms existing state-of-the-art baseline methods and produces stable decisions across scenarios. Furthermore, our method exhibits superior generalization and robustness in cross-dataset detection and adversarial sample attacks, enhancing its reliability in practical applications. We also provide a case study with comparative analysis to demonstrate the practical value of our method.